# HAAP: Vision-context Hierarchical Attention Autoregressive with Adaptive Permutation for Scene Text Recognition

### v1

@@ Abstract
Internal Language Model (LM)-based methods use permutation language modeling (PLM) to solve the error correction caused by conditional independence in external LM-based methods. However, random permutations of human interference cause fit oscillations in the model training, and Iterative Refinement (IR) operation to improve multimodal information decoupling also introduces additional overhead. To address these issues, this paper proposes the Hierarchical Attention autoregressive Model with Adaptive Permutation (HAAP) to enhance the location-context-image interaction capability, improving autoregressive generalization with internal LM. First, we propose Implicit Permutation Neurons (IPN) to generate adaptive attention masks to dynamically exploit token dependencies. The adaptive masks increase the diversity of training data and prevent model dependency on a specific order. It reduces the training overhead of PLM while avoiding training fit oscillations. Second, we develop Cross-modal Hierarchical Attention mechanism (CHA) to couple context and image features. This processing establishes rich positional semantic dependencies between context and image while avoiding IR. Extensive experimental results show the proposed HAAP achieves state-of-the-art (SOTA) performance in terms of accuracy, complexity, and latency on several datasets.
@@ I Introduction
Scene Text Recognition (STR) aims to transcribe the text in an image into a computer-readable text format i.e., to recognize the localized text regions. Scene text with rich information plays a vital role in a range of applications such as visual quizzing, autonomous driving, image retrieval, augmented reality, retail, education, and visually impaired devices [1, 2, 3, 4, 5].
@@
In contrast to Optical Character Recognition (OCR) in documents with more homogeneous text attributes, STR has to deal with irregular text, such as rotation, bending, blurring, or occlusion. In addition, noise and distortion issues make STR challenging. STR is primarily a visual task, involving two different modalities: image and text. Early work still tends to rely on a backbone pre-trained on unimodal data. Learning from data [6, 7, 8, 9, 10, 11, 12, 13] is achieved by combining visual features with semantic priors from word representation models [14] or dictionaries [15] as well as sequence modeling [16]. However, there are inherent limitations in models based on visual information. Visual features alone are not sufficient in cases where the text part is unreadable, e.g., occlusion. Additionally, STR is a One-shot task that has data independence. And, textual semantics can assist visual models in recognizing heterogeneous data.
Zhao et al. verified that language modeling (LM) can effectively focus on text and accurately determine the meaning of words [13]. LM demonstrates the special ability to perceive and understand text in natural images. For that, the combination of LM and visual models for sequence modeling is the current mainstream [8, 9, 10, 11, 12, 13]. It is divided into two categories: external LM-based and internal LM-based. External LM-based schemes essentially accomplish STR by introducing external LMs to assist in correcting visual model reasoning [8, 9, 10, 11]. External LM causes error correction own to the conditional independence of the input. The condition assumes that each pixel or character is independent, regardless of their interactions. However, the reality is that the order and spatial position of the characters about each other affect their meaning and grammatical structure. Algorithms that only consider each character by itself and ignore interactions can lead to prediction errors, especially when the input text has a complex structure and arrangement. Therefore, Permutation Language Modeling (PLM) is employed to improve the generalization of autoregressive (AR) modeling by parsing sequence meaning and syntactic structure [12, 13]. PLM models can be viewed as a collection of AR models with shared architecture and weights. Token dependencies are specified dynamically through the use of an attention mask. Although randomized permutation increases the diversity of data, this strategy is accompanied by training fit oscillations since the random subset has the uncertainty of human interference. We also find that random permutation leads to unstable convergence of the model during training. In addition, if the model only performs well for specific alignments, it leads to a degradation of the generalization performance. Next, recent approaches have used Transformer encoder-decoder based for cross-modal visual and semantic information fusion [8, 9, 10, 11, 12, 13]. However, They need to optimize the predictions through additional Iterative Refinement (IR). Essentially, internal LM is introduced in the decoder can eliminate IR, which is only an additional decoding because the visual and LM are jointly parsed. Considering that noisy inputs still have errors from the output of the visual model, IR is proposed to refine predictions from visual and language cues in external LM-based scenarios. Therefore, recent internal LM-based schemes do not propose an efficient way to couple cross-modal information in the decoder but rely on additional IR.
@@
Figure 2: The basic flow of STR. (a) Visual feature coding and decoding (b) Visual-contextual representation based on external LM. (c) Visual-contextual representation based on internal LM. (d) Internal LM-based visual-contextual representation without Iterative Refinement (IR) (Ours).
This paper proposes the Hierarchical Attention AR Model with Adaptive Permutation (HAAP) to address these issues. The motivation is depicted. First, the use of an adaptive permutation instead of a random permutation scheme increases the diversity of the training data and prevents dependence on a specific order. It reduces the training overhead of the PLM while avoiding training fitting oscillations (as shown in Fig. 1). Second, Multi-head Attention (MHA) is introduced to hierarchically couple cross-modal information. Hierarchical processing improves the adaptation of visual features to semantics and thus enhances end-to-end joint parsing without additional IR (as shown in Fig. 2).
First, inspired by traditional neurons and self-attentive mechanisms, this paper proposes an Implicit Permutation Neurons (IPN) to adapt the ordering between characters in the internal LM. The ordering and spatial location of the characters are considered rather than independently. IPN assigns a set of queries to the original left-to-right permutation to map it linearly into the high-dimensional space. Immediately, two sets of weight matrices are utilized for query weighting and inverse mapping. The process is still able to capture nonlinear relationships in the data with learnable weights. Second, this paper proposes Cross-modal Hierarchical Attention mechanism (CHA) to couple cross-modal information. Contexts embedded in the high dimensional space establish semantic dependencies through self-attention. The context interacts with the image features for the first time to context-image feature alignment. Next, the encoded location is matched with the context and then coupled with the image features again by mutual attention. Hierarchical feature processing drives the interactions of each item in the sequence while capturing the dependencies between locations using an image-guided context-assisted model. In addition, the location-context-image interaction is improved with location inter-correlation guided by adaptive mask and the uniform gradient flow. Therefore, the model avoids additional overhead by eliminating IR and achieves end-to-end parsing.
@@
  • This paper proposes an internal LM-based HAAP to unify end-to-end coupling of location coding, image features, and context to improve the generalization ability of the model in various STR scenes.
  • We propose a novel neuronal structure i.e. IPN to capture the dependencies between characters via adaptive masks efficiently. The model automatically learns the optimal alignment form by capturing the nonlinear relationships during the learning of weights without manual design.
  • We propose hierarchical feature processing CHA to motivate the interaction of each item in the sequence, together with adaptive masks to improve the location-context-image interaction.
@@ II Related work
In this section, we review current deep learning-based STR methods. Recent studies are mainly categorized into two directions, context-independent and context-relevant, based on whether textual semantics are considered or not. We summarize and discuss the research in these two directions separately.
@@ II-A Context-independent STR
Considering that STR is essentially a visual task, the researchers intuitively represent the inputs by constructing a visual model based on deep learning. This is the context-independent scheme i.e., category inference by summarizing visual features. Further, context-independent STR can be categorized into parallel-based reasoning and sequential reasoning. The parallel-based scheme only uses the visual features for prediction without considering the relationship between characters. Its main application is Fully Convolutional Network (FCN) [17] to segment characters at the pixel level. Liao et al [18] recognized characters by grouping segmented pixels into text regions. Wan et al [19] proposed an additional sequential segmentation map that records characters in the correct order. Multi-category classification is unable to accurately construct complete phrases because the output characters are conditionally uncorrelated with each other.
Immediately, the scheme of sequence inference using Recurrent Convolutional Networks (RNN) was proposed to capture the correlation of characters [20]. The most typical is the Connectionist Temporal Classification (CTC) [21] based scheme. The RNN models the sequential sequence of features extracted by a convolutional neural network (CNN) and is trained end-to-end using CTC loss [22]. The RNN receives the current input and the previous hidden state at each time step and outputs the current hidden state. This sequential modeling structure allows the RNN to naturally adapt to the characteristics of sequential data, to parse the syntactic and semantic structures in the text. RNN is also used in a series of Attention-based [23] schemes. Several attempts have explored the enhancement of image encoding and decoding by employing different attentional mechanisms [24, 25, 26]. Character classification by aligning features and character positions, or converting STR into a multi-instance learning problem [7, 27]. Sequence modeling was given a new lease of life with the introduction of transformer [16, 28]. Transformer was then proposed as an alternative to RNN for sequence modeling [29, 30, 31, 32, 33, 34, 35]. However, context-independent methods that rely only on image features for prediction cannot address recognition in low-quality images, especially being less robust to corruption such as occlusion, distortion, or incomplete characters. This limitation motivates the use of language semantics to make recognition models more robust.
@@ II-B Context-relevant STR
Context-dependent STR is a typical cross-modal fusion scenario since the internal interactions between vision and language are constructed. Considering that textual semantics can assist visual models in processing heterogeneous data, the community has recently appeared with a large number of researches to assist recognition by constructing LMs [14, 8, 9, 12, 10, 36, 13, 11]. VisionLAN [8] introduced character masking to guide the visual model to refer to language information in the visual context. MATRN [10] referenced context to enhance visual and semantic features based on spatial coding. STRT [11] designed an iterative text Transformer to predict the probability distribution in a sequence of characters. It is shown that robust models can be constructed by introducing external LMs for explicit modeling. Considering that character sequences are usually modeled in a left-to-right manner, the researcher develops the integrated model to capture twice the amount of information through the bi-directional LM [7, 37, 6, 38]. SRN [7] combined the features of two unidirectional Transformers for prediction, which resulted in twice as expensive both computationally and parametrically. Therefore, ABINet [9] used a novel bidirectional completion network based on bidirectional feature representation. However, the conditional independence of external LM on image features limits its performance and makes error corrections. Subsequently, the internal LM was re-proposed for implicit language modeling to improve visual-contextual interaction. PARSeq [12] used PLM rather than standard AR modeling to learn internal LMs. It supports flexible decoding by using a parameterization that decouples the decoding location from the input context. CLIP4STR [13] introduced an encoder-decoder to process image and text information. The encoder is inherited from CLIP [39], while the decoder uses Transformer.
Internal LR-based schemes mitigate the problem of noisy inputs as well as increase the diversity of the data by dynamically specifying the sequence order through the use of an attention mask. However, the strategy of using randomized permutations introduces training fitting oscillations due to the uncertainty of randomized permutations with human interference. Furthermore, most of these context-dependent STR models use Transformer-based encoder-decoders to implement with IR. However, internal LM-based schemes rely on IR to address noisy inputs in external LM-based models. Thus, the decoding strategy of LM model-based schemes at this stage is inefficient and introduces additional computational effort. This paper proposes to use an adaptive sequence ordering scheme to increase the diversity of the training data and prevent the dependence of the model on a specific order. To further enhance the efficiency of cross-modal information coupling at the decoder side, MHA is developed to Hierarchical coupling information, improving the semantic adaptation of visual features. Ultimately, end-to-end joint parsing is enhanced to avoid additional IR.
@@ III Methodology
In this section, we present in detail the principles and framework of our proposed Hierarchical Attention AR Model with Adaptive Permutation (HAAP), including Implicit Permutation Neurons (IPN) and Cross-modal Hierarchical Attention mechanism (CHA).
@@ III-A Overview
HAAP follows an encoder-decoder architecture as shown in Fig. 3. In the encoding phase, the image and text inputs are represented as a series of patches and semantic tokens. Subsequently, MHA is used for hierarchical image-context interaction and decoding. Specifically, first, the Transformer encoder [28] is utilized to establish the internal feature associations of the image patches. Second, the IPN assigns bi-directional and adaptive permutation masks to the textual context for adjusting sequence order. The contexts embedded into the high-dimensional features are first established by self-attention to establish internal associations. Then MHA is used for context and visual information interaction and position query association. Third, the associated information further interacts with the visual information by utilizing mutual attention. Next, the position query is linearly decoded into contextual output.
Figure 4: Illustration of a ViT layer from Dosovitskiy et al . [28]. $LN$ pertains to layer normalization. $MLP$ represents Multilayer Perceptron.
@@ III-B Encoder
Vision encoder Considering that ViT [28] divides images into patches and uses full connectivity to learn global relationships can better extract global and complex visual information than traditional CNNs [40], HAAP follows the original ViT structure to build the visual encoder, which consists of $L=12$ layers of Transformer encoder. All layers share the same architecture as shown in Fig. 4. The output of the last encoder is subjected to a Layer Normalization (LN). Formally, first, the image of size $H\times W\times C$ (height $H$, width $W$, channels $C$) is reshaped into a series of flat 2D patches $x_{p}\in\mathbb{R}^{N\times(P^{2}-C)}$, where $(P,P)$ is the resolution of each image block and $N=HW/P^{2}$ is the number of blocks generated. The patch $x_{p}$ is flattened and mapped to $D$ dimensions using a trainable linear projection $E\in\mathbb{R}^{({P^{2}}C)\times D}$. The position embedding $E_{pos}\in\mathbb{R}^{(N+1)\times D}$ is added to the projected output $z$, as:
@@
where $x^{i}_{p}$ represents the $i$-th $x_{p}$. $z$ performs feature extraction in alternating MHA and Mult-Layer Perceptron (MLP). A normalized $LN$ is applied initially in each block to mitigate internal covariate bias during training. Additionally, residual concatenation is applied twice in the block to aid gradient propagation in the depth model, as:
@@
where $MLP$ contains two nonlinear layers with GELU activation functions. The $z^{\prime}_{l}$ and $z_{l}$ are the visually encoded information. MHA is the extension of scaled dot-product attention to multiple representation subspaces or heads. In scaled dot-product attention, the similarity scores are computed using dot-product computation of the similarity between two $d_{k}$-dimensional vectors query $q$, and key $k$. Attention scores are transformed by the $d_{v}$-dimensional vector values $v$. The scaled dot product attention is defined as:
$Attn(q,k,v)=softmax(\frac{qk^{T}}{\sqrt{d_{k}}})v,$ (4)
further,
$MHA(q,k,v)=Concat(head_{1},\cdots head_{h})W^{o},$ (5)
where
$head_{i}=Attn(qW^{q}_{i},kW^{k}_{i},vW^{v}_{i}),$ (6)
where the computational cost of MHA is practically constant due to the dimension of the vector $d_{head}=d_{model}/h$, where $h$ is the number of heads. The $W^{q},W^{k},W^{v}\in\mathbb{R}^{d_{model}\times d_{head}}$ are obtained by connecting the heads and multiplying with the output projection matrix $W^{o}\in\mathbb{R}^{d_{model}\times d_{head}}$.

### v2

@@ Abstract
Scene Text Recognition (STR) is challenging in extracting effective character representations from visual data when text is unreadable. Permutation language modeling (PLM) is introduced to refine character predictions by jointly capturing contextual and visual information. However, in PLM, the use of random permutations causes training fit oscillation, and the iterative refinement (IR) operation also introduces additional overhead. To address these issues, this paper proposes the Hierarchical Attention autoregressive Model with Adaptive Permutation (HAAP) to enhance position-context-image interaction capability, improving autoregressive LM generalization. First, we propose Implicit Permutation Neurons (IPN) to generate adaptive attention masks that dynamically exploit token dependencies, enhancing the correlation between visual information and context. Adaptive correlation representation helps the model avoid training fit oscillation. Second, the Cross-modal Hierarchical Attention mechanism (CHA) is introduced to capture the dependencies among position queries, contextual semantics and visual information. CHA enables position tokens to aggregate global semantic information, avoiding the need for IR. Extensive experimental results show that the proposed HAAP achieves state-of-the-art (SOTA) performance in terms of accuracy, complexity, and latency on several datasets.
@@ I Introduction
Scene Text Recognition (STR) aims to transcribe text in an image to a computer-readable text format, i.e., to recognize the localized text regions. Scene text with rich information plays a vital role in a range of applications such as visual quizzing, autonomous driving, image retrieval, augmented reality, retail, education, and visually impaired devices [40, 42, 26, 5, 23].
@@
In contrast to Optical Character Recognition (OCR) in documents with more homogeneous text attributes, STR poses significant challenges due to irregular text, including rotated, curved, blurred, and occluded characters. STR is primarily a vision-based task that involves two different modalities: image and text. Early works use a model pre-trained on image data and directly classify characters according to visual information of the text region without incorporating context, i.e., the meaning of text [36, 55]. However, visual information alone is insufficient in cases where the text is unreadable, e.g., occluded, blurred, or with various fonts and spatial layouts [4]. By language modeling (LM), text is introduced to aid recognition as text carries rich contextual information. Modeling the dependencies between the visual information and context and using autoregressive sequence modeling to predict character tokens [47, 11, 29, 49, 4, 59, 51].
Current LM-based approaches can be divided into two categories: external LM-based [47, 11, 29, 49] and internal LM-based [4, 59, 51]. The external LM-based approach introduces an independent LM to correct the token prediction of the visual model by explicitly modeling the context. In contrast, the internal LM-based approach embeds LM in the visual model to exploit joint visual information and context. External LM is used as a spell checker to rectify the visual prediction via unidirectional sequence modeling. The conditional independence of the LM from the input may cause it to correct the predictions erroneously [11, 49]. Therefore, the internal LM-based approach applies Permutation Language Modeling (PLM) to enhance the correlation between the input image and context, improving the generalization of sequence modeling [4, 59]. It can be seen as an ensemble of autoregressive models with shared architecture and weights [4]. Using attention masks to dynamically specify token dependencies, PLM allows the decoder to perform sequence modeling of characters in arbitrary orders without relying on specific sequence order [59]. Compared to unidirectional modeling, such as left-to-right (LTR) and right-to-left (RTL), the model can capture more information about token dependencies.
However, there are still two problems: 1:) PLM produces a training fit oscillation problem due to the introduction of randomized permutation. Randomized permutation brings stochastic token dependencies that cause the model to have an unstable fitting process in decoding, which is defined as training fit oscillation in this paper. This makes decoding inefficient and requires multi-training to determine the best model. 2:) Iterative refinement (IR) was initially proposed in external LM-based approaches [47, 11, 29, 49] and then in internal LM-based ones [4, 59, 51] to aggregate global information from visual and language models to refine prediction. However, internal LM-based models have adopted an attention-based encoder-decoder architecture that integrates visual and language modeling. It means that contextual semantics can be refined from visual information in the decoder instead of using IR, which is inefficient and increases decoding time and complexity.
@@
Figure 2: The basic flow of STR. (a) Visual feature coding and decoding (b) Joint visual and context representation based on external LM. (c) Joint visual and context representation based on internal LM. (d) Internal LM-based visual-context representation without Iterative Refinement (IR) (Ours).
To avoid the training fit oscillation and the use of IR, this paper proposes the Hierarchical Attention autoregressive model with Adaptive Permutation (HAAP) to improve the efficiency of cross-modal information decoding. In particular, first, we propose Implicit Permutation Neurons (IPN) to establish an adaptive attention mask in decoding by jointly considering the visual information and context (as illustrated in Fig. 1). IPN adaptively exploits token dependencies, enhancing the input-output correlations within the model and preventing the training fit oscillation. Second, we develop the Cross-modal Hierarchical Attention mechanism (CHA) to integrate cross-modal information by establishing the dependencies among the position queries, context and the visual information using the adaptive attention mask provided by IPN. The contextual semantics are refined through a hidden layer, enabling each position token to be predicted based on global contextual semantics, rather than solely on preceding tokens. CHA significantly improves decoding efficiency and avoids the need for IR (as shown in Fig. 2).
@@
  • This paper proposes an internal LM-based HAAP to end-to-end coupling of position, image, and context to improve the generalization ability of the model in various STR scenes.
  • We propose a novel neuronal structure, i.e., IPN, to efficiently capture the token dependencies via an adaptive attention mask. The model automatically learns contextual semantic representations.
  • A hierarchical feature processing CHA is introduced to establish the position-context-image interaction. Enhancing the correlation between visual information and context and expanding the semantic receptive field of each position token.
@@ II Related work
In this section, we review current deep learning-based STR methods. Recent studies are mainly categorized as context-independent and context-relevant. We summarize and discuss the research in these two directions separately.
@@ II-A Context-independent STR
Since STR is a visual task, the researchers intuitively represent the input using a visual model and perform the class prediction from visual features. The parallel prediction-based methods directly use Fully Convolutional Network (FCN) [25] to segment characters pixel by pixel without considering the relationship between characters. Liao et al [27] recognized characters by grouping segmented pixels into text regions. Wan et al [45] identified characters in the correct order using an additional segmentation map. Multi-class prediction cannot accurately construct a complete text instance because the output characters are conditionally uncorrelated.
Immediately, sequence modeling-based methods use Recurrent Convolutional Networks (RNN) to capture the correlation of characters [24]. The most typical is the Connectionist Temporal Classification (CTC)-based scheme [12]. The RNN models the feature sequence extracted by a convolutional neural network (CNN) and is end-to-end trained using CTC loss [35]. The RNN receives the current input and the previous hidden state at each time step and outputs the current representation. This sequence modeling structure allows the RNN to naturally adapt to the characteristics of sequential data, to capture the linguistic rules and semantics in the text. RNN is also used in a series of Attention-based [3] schemes. Several attempts have explored the enhancement of image encoding and decoding by employing different attention mechanisms [6, 19, 20]. Character classification by aligning features and character positions, or converting STR into a multi-instance learning problem [55, 14]. Sequence modeling-based methods have made great progress with the introduction of the Transformer [43, 10]. Transformer was then proposed as an alternative to RNN for sequence modeling [1, 2, 48, 21, 52, 50, 9]. However, context-independent methods that rely only on image features for prediction cannot address recognition in low-quality images, especially being less robust to corruption such as occlusion, distortion, or incomplete characters. This limitation motivates using context to make recognition models more robust.
@@ II-B Context-relevant STR
Context-dependent STR is a typical cross-modal fusion process since the internal interactions between vision and context are constructed. Considering that contextual semantics can assist visual models in processing heterogeneous data, the community has recently appeared with a large number of research studies using LM to assist recognition [32, 47, 11, 4, 29, 57, 59, 49]. VisionLAN [47] introduced character masking to guide the visual model to refer to text information in the visual context. MATRN [29] referenced context to enhance visual semantics based on spatial coding. STRT [49] designed an iterative text Transformer to predict the probability distribution in character sequences. It is shown that robust recognition models can be constructed by introducing external LMs for explicit modeling. Considering that character sequences are usually modeled in a left-to-right manner, the researcher develops the integrated model to capture twice the amount of information through the bi-directional LM [55, 36]. SRN [55] combined the features of two unidirectional Transformers for prediction, which resulted in twice as expensive both computationally and parametrically. Therefore, ABINet [11] used a novel bidirectional completion network based on bidirectional feature representation. STRT [49] used LM to capture the linguistic rule and refine the position token prediction. However, the conditional independence of external LM from the input image limits its performance and makes the model erroneously rectify the correct predictions. Subsequently, the internal LM was proposed for implicit language modeling to improve image-context interaction. PARSeq [4] used PLM to enhance the generalization of internal LM. It supports flexible decoding by dynamically specifying the dependence of position tokens and then decoding the position tokens from the input. CLIP4STR [59] introduced CLIP [33] and used PLM to construct an encoder-decoder framework to integrate image and text information.
PLM-based methods mitigate the problem of noisy inputs and increase the diversity of the data by dynamically specifying the sequence order. However, the use of randomized permutations introduces the problem of training fit oscillation. Furthermore, most internal LM-based models employ an attention-based encoder-decoder architecture to integrate visual information with contextual semantics. Nevertheless, due to inefficient decoding, these models typically adopt IR to refine prediction, at the cost of increased computational complexity. Randomized token dependencies make an insufficient contextual semantic representation. Since each position token prediction only relies on the semantics preceding the tokens, the model needs to use IR further to aggregate global contextual semantics. This paper proposes to adopt an adaptive sequence modeling strategy that promotes training data diversity and effectively captures the interdependence between visual and contextual information. To further enhance the efficiency of cross-modal information decoding, the Multi-head attention mechanism (MHA) is used to hierarchically model relationships among position queries, contextual semantics, and visual information. The correlation between position and context is enhanced, enabling each token to access global contextual semantics and avoiding the need for IR.
@@ III Methodology
In this section, we present in detail the principles and framework of our proposed Hierarchical Attention autoregressive Model with Adaptive Permutation (HAAP), including Implicit Permutation Neurons (IPN) and Cross-modal Hierarchical Attention mechanism (CHA).
@@ III-A Overview
HAAP follows an encoder-decoder architecture as shown in Fig. 3. In the encoding phase, the image and text inputs are represented as a series of patches and context tokens. Subsequently, MHA is used for hierarchical image-context interaction and decoding. Specifically, first, the Transformer encoder [10] is used to establish the internal feature representation and correlation of the image patches. Second, the IPN assigns bi-directional and adaptive attention masks to the context for adjusting the order of autoregressive sequence modeling. The first attention stage establishes the dependencies between the visual information and context with attention masks, producing the contextual semantics. The second attention stage encodes the contextual semantics and position through MHA into position queries. Finally, guide with visual information, position queries autoregressive decode the position token per iteration.
Figure 4: Illustration of a ViT layer from Dosovitskiy et al. [10]. LNLN pertains to layer normalization. MLPMLP represents Multilayer Perceptron.
@@ III-B Encoder
Vision encoder As shown in Fig. 4, HAAP uses $L=12$ layers of the Transformer encoder [10] to build the visual encoder. The output of the last encoder is processed by layer normalization $LN$. First, the image of size $H\times W\times C$ (height $H$, width $W$, channels $C$) is reshaped into a series of flat 2D patches $x_{p}\in\mathbb{R}^{N\times(P^{2}-C)}$, where $(P,P)$ is the resolution of each image block and $N=HW/P^{2}$ is the number of blocks generated. The patch $x_{p}$ is flattened and mapped to $D$ dimensions using a trainable linear projection $E\in\mathbb{R}^{({P^{2}}C)\times D}$. The position embedding $E_{pos}\in\mathbb{R}^{(N+1)\times D}$ is added to the projected output $z$, as:
@@
where $x^{i}_{p}$ represents the $i$-th $x_{p}$. $z$ performs feature extraction in alternating MHA and Mult-Layer Perceptron (MLP). $LN$ is applied initially in each block to mitigate internal covariate shift during training. Additionally, residual concatenation is applied twice in the block to aid gradient propagation in the depth model, as:
@@
where $MLP$ contains two nonlinear layers with GELU activation functions. The $z^{\prime}_{l}$ and $z_{l}$ are encoded visual information. MHA is the extension of scaled dot-product attention to multiple representation subspaces or heads [4].